f(x)= 1, for x≥1
-4-x, for x<1

Answer by stanbon(75887)

f(x)= 1, for x≥1
f(x) = -4-x, for x<1
-----
You probably need to graph this function.
----
Graph the line y = -x-4 from x > -oo to x = 1
Graph the horizontal line y = 1 from x = 1 to +oo
-----------------
Cheers,
Stan H.
------------------